Circular No. 7442

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cq IN UGC 10354 W. D. Li, University of California at Berkeley, on behalf of the Lick

Observatory Supernova Search (cf. IAUC 6627, 7126), reports the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on June 23.3 UT, with confirmation images taken taken on June 24.3 (limiting mag ∼ 20.5) and 25.3 (limiting mag ∼ 21.5), all showing the new object at mag ∼ 19.3. SN 2000cq — which also showed a hint of presence in images taken on May 30, June 2, 11, 15, and 20 — is located at α = 16^{h}21^{m}21^{s}.34, δ = +40^{o}48′09′′.7 (equinox 2000.0), which is 7′′.8 west and 28′′.6 south of the nucleus of UGC 10354. KAIT images of the same field taken on May 3.3 and 9.3 (all with limiting mag ∼ 20.0) showed nothing at the position of the supernova, and there is no indication of SN 2000cq on the second Palomar Sky Survey blue (1989) and red (1993) images (limiting mag ∼ 22.0).

SUPERNOVA 2000cp IN PGC 57064 S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that a spectrum of SN 2000cp (cf. IAUC 7441), obtained by P. Berlind on June 22.28 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova at maximum light. Narrow superimposed Hα emission yields a recession velocity of 10 260 km/s for the host galaxy. The expansion velocities of the Ca II H and K and the Si II (rest 635.5 nm) features are 16 400 and 11 400 km/s, respectively. The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 1 ± 3 days before maximum light.

AX J0049.4–7323 M. Ueno, J. Yokogawa, K. Imanishi, and K. Koyama, Kyoto Univer-

sity, communicate: “An ASCA observation made during Apr. 11–17 revealed coherent pulsations of period 755.5 ± 0.6 s from a new source in the Small Magellanic Cloud, designated AX J0049.4–7323 and located at α = 0^{h}49^{m}25^{s}.2, δ = −73^{o}23′17′′ (equinox 2000.0; error radius 1′.5). The spectrum is characterized by a flat power-law function with photon index 0.7 and x-ray flux 1.1 × 10^{−12} erg cm^{−2} s^{−1} (0.7–10 keV). The Be/x-ray binary RX J0049.7–7323 (Haberl and Sasaki 2000, http://xxx.lanl.gov/abs/ astro-ph/?0005226) is located within the ASCA error region.”

2000 June 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green